# Breaking Annotation Barriers: Generalized Video Quality Assessment via Ranking-based Self-Supervision

### v2

@@ 4.1 Model Structure
Methods Training data Label SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
Traning-free / Opinion-unaware VQA Methods
NIQE [38] None 0.240 0.247 0.473 0.496 0.354 0.413 0.048 0.002 0.163 0.114 0.183 0.154
IL-NIQE [72] 0.200 0.168 0.340 0.303 -0.081 -0.040 0.079 -0.008 0.056 0.006 0.083 0.036
VIIDEO [39] 0.077 -0.199 0.157 -0.257 0.276 0.244 0.114 0.078 0.082 0.019 0.110 0.010
STEM [20] 0.103 0.111 0.498 0.492 0.288 0.317 0.184 0.097 0.123 0.104 0.172 0.147
TPQI [27] NA NA NA NA NA NA NA NA NA NA NA NA
BUONA-VISTA [63] NA NA NA NA NA NA NA NA NA NA NA NA
[…]
MinimalisticVQA(VII) [52] LSVQ [70] 0.596 0.682 0.681 0.733 0.061 0.130 0.275 0.338 0.604 0.659 0.490 0.551
MinimalisticVQA(IX) [52] LSVQ [70] 0.686 0.746 0.797 0.816 0.301 0.388 0.459 0.502 0.615 0.661 0.574 0.622
FAST-VQA [61] LSVQ [70] 0.631 0.677 0.725 0.747 0.326 0.415 0.327 0.363 0.518 0.526 0.486 0.512
DOVER [64] LSVQ [70] [40] 0.647 0.728 0.694 0.747 0.360 0.465 0.368 0.418 0.559 0.593 0.519 0.569
Q-Align [65] fused [17, 11, 28, 70, 40] 0.611 0.681 0.756 0.798 0.329 0.342 0.414 0.497 0.613 0.655 0.555 0.606
[…]
Hard Ranking PLVD-P1 (500k) 0.705 0.742 0.794 0.801 0.496 0.550 0.492 0.542 0.640 0.670 0.613 0.648
Soft Ranking (Stage 1) PLVD-P1 (500k) 0.697 0.752 0.799 0.829 0.481 0.525 0.552 0.614 0.690 0.725 0.650 0.693
Soft Ranking (Stage 2) PLVD-P1/ P2 (600k) 0.717 0.763 0.810 0.834 0.515 0.611 0.633 0.696 0.749 0.764 0.704 0.741
Soft Ranking (Stage 3) PLVD-P1/ P2/ P3 (700k) 0.703 0.761 0.792 0.823 0.559 0.644 0.670 0.728 0.754 0.770 0.716 0.752
4.1.1 Training and Inference Pipelines
Training. LMM-PVQA is optimized via standard supervised fine-tuning (SFT). To enhance training efficiency, the LLM parameters are kept frozen to preserve their intrinsic reasoning capabilities. Similarly, the pretrained SlowFast motion encoder remains frozen due to its proven effectiveness in capturing temporal representations [52]. Trainable parameters are confined to the image encoder $\mathcal{F_{I}}$ and feature projection layers $\mathcal{P}_{I}$ and $\mathcal{P}_{M}$, thus enabling domain-specific visual feature adaptation.
Furthermore, we propose an iterative self-improvement training strategy. The training process consists of three stages. First, the base model is trained on the PLVD-Part1 dataset. Next, the trained model acts as an enhanced judge, annotating video pairs in the PLVD-Part2 dataset alongside existing judges. The combined PLVD-Part1 and PLVD-Part2 datasets are then used to train a second-stage model. This process is repeated for a third-stage model. As demonstrated in Section 5.2, this iterative strategy progressively enhances out-of-distribution performance.
Inference. Given that the model primarily predicts relative quality rankings between video pairs, a practical conversion to absolute quality scores is required. To achieve this, we adopt an adaptive soft comparison method proposed in  [77], which first computes a soft probability matrix across ranking categories by comparing the test video against anchor videos, followed by maximum a posteriori (MAP) [54] estimation under Thurstone’s Case V model [53] to obtain calibrated quality scores.
For anchor selection, the range of quality scores of PLVD-VQA is partitioned into five quality intervals. Within each interval, the video exhibiting the lowest inter-model score variance is selected as the anchor, ensuring stable reference points for score derivation. We list the detailed calculation procedure in the supplement file.
5 Experiments
5.1 Experimental Setups
Implementation Details For each stage of training, we initialize the image encoder, image projector, and LLM with LLaVA-ov-chat (7B) weights [24] and the motion encoder with SlowFast pre-trained weights. The model is trained for one epoch on four NVIDIA A800 GPUs with a total batch size of $4$ and a learning rate of $1$e-$5$. Inference is performed on two NVIDIA RTX 3090 GPUs.
Benchmark Datasets The proposed methods is evaluated on ten VQA benchmarks, categorized into in-domain and out-of-distribution (OOD) groups to systematically assess their generalization. The in-domain evaluation benchmarks include LSVQ Test [70], LSVQ 1080p [70], KoNViD-1k [16], LIVE-VQC [50], and YouTube-UGC [60], all of which contain UGC videos. The OOD datasets consist of LIVE-YT-Gaming [71], CGVDS [47], LIVE-YT-HFR [33], Waterloo-IVC-4K [26], and KVQ [32]. Specifically, LIVE-YT-Gaming and CGVDS contain gaming videos, which exhibit significant differences in content compared to natural UGC videos. LIVE-YT-HFR includes videos with varying frame rates, focusing on temporal-related distortions. Waterloo-IVC-4K and KVQ primarily evaluate compression artifacts, where Waterloo-IVC-4K contains compressed high-resolution videos at different spatial resolutions, while KVQ comprises compressed UGC videos processed with various video enhancement techniques. We introduce details of these datasets in the supplement file.
Evaluation Criteria We adopt two widely used criteria to evaluate the performance of VQA models: Spearman Rank Correlation (SRCC) and Pearson Linear Correlation (PLCC), which indicate the prediction monotonicity and prediction linearity, respectively.
Competing Methods We compare our model with six opinion-unaware BVQA methods, including NIQE [38], IL-NIQE [72], VIIDEO [39], STEM [20], TPQI [27], and BUONA-VISTA [63], as well as five supervised VQA methods: Minimalistic-VQA (VII) [52], Minimalistic-VQA (IX) [52], FAST-VQA [61], DOVER [64], and Q-Align [65]. These five supervised VQA models serve as strong baselines, as they are used to pseudo-label video quality and can be regarded as teacher models. The supervised models are trained on the LSVQ training set.
Table 3: Supervised performance comparison of our model against competitive methods. The experimental results on KoNViD-1k, YouTube-UGC, and LIVE-VQC are obtained using 10-fold cross-validation. “Base” denotes training LMM-PVQA without self-supervised pretraining
[…]
TLVQM [22] 0.772 / 0.774 0.589 / 0.616 0.732 / 0.724 0.670 / 0.691 0.685 / 0.692
VIDEAL [55] 0.794 / 0.783 0.545 / 0.554 0.751 / 0.741 0.630 /0.640 0.687 / 0.709
VSFA [25] 0.801 / 0.796 0.675 / 0.704 0.794 / 0.799 0.718 / 0.771 0.787 / 0.789
PatchVQ [70] 0.827 / 0.828 0.711 / 0.739 0.791 / 0.786 0.827 / 0.837 NA / NA
SimpleVQA [51] 0.867 / 0.861 0.764 / 0.803 0.856 / 0.860 0.845 / 0.859 0.847 / 0.856
FAST-VQA [61] 0.880 / 0.880 0.781 / 0.813 0.891 / 0.892 0.849 / 0.865 0.855 / 0.852
DOVER [64] 0.878 / 0.866 0.782 / 0.813 0.908 / 0.910 0.860 / 0.875 0.841 / 0.851
MinimalisticVQA [52] 0.881 / 0.879 0.781 / 0.820 0.889 / 0.890 0.842 / 0.854 0.890 / 0.891
Soft Ranking (Base) 0.880 / 0.864 0.790 / 0.814 0.896 / 0.877 0.840 / 0.853 0.858 / 0.848
Soft Ranking (Stage 1) 0.907 / 0.904 0.832 / 0.857 0.911 / 0.908 0.884 / 0.894 0.910 / 0.911
Table 4: Experimental results of ablation study. “Spatial” and “Motion” denote the spatial branch and the temporal branch of the visual feature extractor. In Stage 1, if PLVD-DS (Part1) is not specified, the model is trained solely on PLVD-VQA (Part1). In Stage 2, “w/o label refinement” and “w/ label refinement” indicate whether the trained model is used to label the data in PLVD-Part2
[…]
Spatial Motion PLVD-DS w/o label refinement w/ label refinement SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
0.872 0.862 0.787 0.814 0.868 0.864 0.783 0.819 0.834 0.839 0.843 0.846
0.883 0.872 0.804 0.827 0.883 0.876 0.799 0.830 0.842 0.843 0.855 0.857
0.886 0.880 0.803 0.830 0.891 0.888 0.797 0.832 0.845 0.849 0.858 0.863
[…]
0.887 0.880 0.802 0.830 0.893 0.892 0.798 0.833 0.856 0.859 0.859 0.864
Stage 1 Stage 2 LIVE-YT-Gaming CGVDS LIVE-YT-HFR Waterloo-IVC-4K KVQ Overall
Spatial Motion PLVD-DS w/o label refinement w/ label refinement SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
0.678 0.735 0.754 0.800 0.334 0.413 0.404 0.459 0.616 0.656 0.561 0.610
0.688 0.736 0.769 0.808 0.446 0.477 0.425 0.470 0.619 0.662 0.579 0.622
0.697 0.752 0.799 0.829 0.481 0.525 0.552 0.614 0.690 0.725 0.650 0.693
[…]
0.717 0.763 0.810 0.834 0.515 0.611 0.633 0.696 0.749 0.764 0.704 0.741
5.2 Performance Analysis
The performance of all compared methods and our proposed models is summarized in Table 4.1, with detailed analysis conducted from the following perspectives:
Hard Ranking v.s. Soft Ranking. The model trained with soft ranking marginally outperforms its hard-ranking counterpart on in-domain VQA benchmarks and CGVDS. Moreover, it achieves significantly better performance on Waterloo-IVC-4K and KVQ, while exhibiting slightly inferior results on LIVE-YT-Gaming and LIVE-YT-HFR. Overall, the soft ranking strategy incorporates “quality distance” information into the training process, enabling the model to easily learn quality comparison capabilities and achieve better performance.
Iterative Self-Improvement Training Strategy. Consistent performance improvements are observed across both in-domain and OOD datasets as the model advances through training stages. This empirical evidence demonstrates that our iterative training strategy effectively enhances model capability through progressive self-learning. Notably, substantial gains are achieved on challenging VQA benchmarks where existing models struggle: the LIVE-YT-HFR, Waterloo-IVC-4K, and KVQ datasets exhibit relative improvements of $16.2\%$, $21.4\%$, $9.3\%$ in SRCC respectively, after three-stage refinement.
In-Domain Performance Comparison. We observe that all proposed model variants outperform the five competing VQA models, which are used for label annotation in Section 3.2.1, confirming our self-supervised approach successfully distills ensemble knowledge from its teacher models on unlabeled data. However, our models are slightly outperformed by FAST-VQA and DOVER on LIVE-VQC, which exhibits more complex temporal distortions compared to other datasets [52]. We attribute this performance gap to architectural differences: FAST-VQA and DOVER fine-tune a 3D DNN (i.e., VideoSwin) specifically optimized for spatiotemporal feature extraction, whereas our framework handles temporal information using a pre-trained SlowFast model.
OOD Performance Comparison. Our models significantly outperform competing approaches on the OOD benchmarks, achieving a $24.7\%$ improvement in SRCC for overall OOD evaluation compared to the competing approaches. We attribute the strong OOD performance to two key factors. First, our training datasets include compression distortions (H.264 and H.265), and the synthetic quality ranking labels provide useful supervision for assessing compression-related video quality for these datasets. Second, our iterative self-improvement training strategy enables progressive adaptation to unseen distortion types, such as frame rate inconsistencies in LIVE-YT-HFR, AVS/VP9 compression artifacts in Waterloo-IVC-4K, and enhancement-induced distortions in KVQ—none of which are present in the training data. These results demonstrate the effectiveness of the proposed self-supervised VQA framework in real-world quality assessment scenarios.
5.3 Supervised Performance Comparison
We conduct supervised fine-tuning of our self-supervised trained model on the in-domain VQA benchmarks, as shown in Table 3. The results reveal that: (1) Our fine-tuned model surpasses all competing VQA baselines with 3.5% higher SRCC, and (2) it achieves $6.4\%$ improvement over the baseline trained without self-supervised initialization. These results confirm that self-supervised representation learning critically enhances downstream fine-tuning effectiveness.
Furthermore, we compare the prediction accuracy of pairwise comparisons in LMM-PVQA with four open-source LMMs that have demonstrated strong performance in video understanding. As shown in Table 6, LMM-PVQA significantly outperforms these advanced LMMs, providing highly accurate quantitative predictions. In contrast, other models primarily rely on high-level instruction-tuning datasets, resulting in suboptimal accuracy in quality prediction. This suggests that existing LMMs exhibit limited quality perception capabilities, highlighting the advantages of our proposed approach in video quality assessment.
5.4 Ablation Study
We conduct the ablation experiments to validate the effectiveness of LMM-PVQA, with results shown in Table 4.
Motion Encoder. Incorporating the motion encoder for temporal distortion representation yields consistent improvements across all benchmarks, particularly achieving $33.5\%$ SRCC improvement on LIVE-YT-HFR—the benchmark specifically designed for high frame-rate distortion analysis. This verifies that explicit motion modeling is essential for capturing temporal-related distortions.
Synthetic Distortion Data. The incorporation of the PLVD-DS dataset also demonstrates consistent performance improvements across all benchmarks, yielding marginal gains on in-domain benchmarks while achieving substantial enhancements on OOD benchmarks. It can be attributed that the PLVD-VQA dataset have already equipped the model with quality assessment capabilities for in-domain content, whereas the PLVD-DS dataset help mitigate the domain gap on OOD benchmarks by simulating common artificial degradations (e.g., compression).
Iterative Self-Improvement Training Strategy. Since our iterative self-improvement training strategy introduces new training samples, it is critical to verify that performance gains stemming from the strategy itself rather than additional data. To isolate this effect, we train a Stage 2 model using the same base training set (PLVD-Part1/Part2), where videos in PLVD-Part2 are labeled by five VQA judges, without incorporating the Stage 1 model as the new judge. Experimental results confirm that merely increasing training data (without iterative refinement) fails to surpass the Stage 1 model’s performance, thereby validating the effectiveness of our strategy in enabling self-improvement through iterative optimization.
@@ 6 Conclusion
We propose a self-supervised learning framework for VQA to mitigate dependence on human-annotated datasets. Our approach constructs a large-scale video pair dataset labeled via established VQA models and synthetic distortion simulations, then adopts a learning-to-rank paradigm to learn quality assessment capabilities. By integrating an iterative self-improvement training strategy, our model progressively enhances its evaluation performance through self-learning. Our method achieves state-of-the-art zero-shot performance on both in-domain and OOD benchmarks, demonstrating its effectiveness.
Limitations. Our method is inherently designed for pairwise quality comparison. To obtain absolute quality scores for a single video, it must be compared against multiple anchor videos (five in this study), which significantly increases inference time.
Future Work. We plan to explore more automated video pair annotation strategies, such as leveraging expert-domain VQA models (e.g., VMAF for video compression), utilizing LMMs with carefully designed prompt engineering, and employing text-to-video generation algorithms to synthesize videos of varying quality through specified prompts, thus substantially expanding the diversity and scale of training data. Furthermore, we intend to extend our framework to incorporate additional modalities, such as images and audio, to develop a more generalizable quality assessment model.
Supplementary Material
Figure 4: Examples videos of different categories in our large dataset.
Appendix A More Details of Our PLVD Database
@@ A.1 Analysis of the Collected Videos
As shown in Fig. 5, our dataset is collected from multiple popular social media platforms with relatively uniform sampling, comprising $20\%$ from Bilibili, $20\%$ from Youku, $25\%$ from YouTube, and $35\%$ from TikTok. Notably, our dataset covers a diverse range of content categories, exceeding twenty in total. In addition to common categories such as lifestyle, food, and animals, it also includes specialized categories such as gaming, AI-generated content, and high-resolution content. To illustrate the diversity of our dataset, we present a variety of video samples in Fig. 4, showcasing the broad range of content available in our large-scale video quality assessment (VQA) dataset. Unlike existing datasets, which often focus on specific formats, our dataset encompasses a wider variety of formats, including both landscape and portrait orientations, as well as various resolutions. This diversity enhances the comprehensiveness of our dataset, making it more suitable for evaluating video quality across a wide kinds of scenarios.

### v3

@@ 4.3 Training Strategy
Methods SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
State-of-the-art VQA Methods - teachers
MinimalisticVQA(VII) (Sun et al., 2024) 0.596 0.682 0.681 0.733 0.061 0.130 0.275 0.338 0.604 0.659 0.490 0.551
MinimalisticVQA(IX) (Sun et al., 2024) 0.686 0.746 0.797 0.816 0.301 0.388 0.459 0.502 0.615 0.661 0.574 0.622
FAST-VQA (Wu et al., 2022) 0.631 0.677 0.725 0.747 0.326 0.415 0.327 0.363 0.518 0.526 0.486 0.512
DOVER (Wu et al., 2023a) 0.647 0.728 0.694 0.747 0.360 0.465 0.368 0.418 0.559 0.593 0.519 0.569
Q-Align (Wu et al., 2023b) 0.611 0.681 0.756 0.798 0.329 0.342 0.414 0.497 0.613 0.655 0.555 0.606
State-of-the-art VQA Methods - others
VQA2 (Jia et al., 2024) 0.613 0.698 0.656 0.741 0.332 0.413 0.415 0.474 0.678 0.689 0.583 0.623
VQAThinker (Cao et al., 2025) 0.767 0.806 0.856 0.845 0.528 0.610 0.573 0.624 0.586 0.626 0.615 0.658
Our Weak-to-Strong VQA Methods
(I): Ensembling homogeneous teachers 0.688 0.756 0.769 0.808 0.456 0.497 0.455 0.502 0.649 0.682 0.602 0.643
(II): (I) + Integrating heterogeneous teachers 0.697 0.752 0.799 0.829 0.481 0.525 0.552 0.614 0.690 0.725 0.650 0.693
(III): (II) + Confidence loss 0.708 0.763 0.796 0.829 0.523 0.606 0.579 0.643 0.713 0.742 0.672 0.717
(IV): (III) + Iterative stage W2S training 0.711 0.770 0.807 0.831 0.606 0.678 0.657 0.737 0.759 0.782 0.722 0.765
(V): (IV) + Iterative stage W2S training 0.723 0.776 0.799 0.828 0.683 0.749 0.698 0.758 0.772 0.807 0.745 0.789
4.4 Experimental Results
We present the experimental results in Table 1, highlighting five progressively enhanced models of our method: models (I)–(III) incrementally add components in Stage 1, while model (IV) and model (V) introduce iterative training in Stage 2 and Stage 3, respectively. We analyze them from the following aspects:
Ensembling Homogeneous Teachers. Compared with single-teacher supervision, we find that ensembling multiple teachers yields stronger student models that outperform all individual teachers as well as their corresponding students. This result further highlights the weak-to-strong effect in VQA and shows that improving the quality of teacher supervision amplifies this effect, consistent with prior findings.
Integrating Heterogeneous Teachers. We incorporate synthetic distortion simulators as specialized VQA models to extend the capability of the teacher ensemble. With synthetic distortion pairs, the student model achieves consistent improvements across all benchmarks, yielding marginal gains on in-domain datasets and substantial enhancements on OOD benchmarks. These results demonstrate that incorporating diverse VQA models as teachers enables joint supervision that consistently fosters more generalizable quality assessment.
Confidence Loss. Incorporating $\mathcal{L}_{\text{conf}}$ yields clear gains on OOD datasets. This indicates that confidence loss mitigates the adverse impact of noisy weak labels and enables the student to reinforce its own reliable predictions.
Iterative W2S Training. We observe consistent improvements across both in-domain and OOD datasets as the student progresses through three iterative training stages. This provides strong empirical evidence that our iterative weak-to-strong strategy enhances model capacity through progressive self-teaching. Notably, substantial gains are achieved on challenging benchmarks where existing models struggle: after three iterations, relative SRCC improvements of $30.59\%$, $20.55\%$, and $8.27\%$ are obtained on LIVE-YT-HFR, Waterloo-IVC-4K, and KVQ, respectively.
Comparison with SOTAs. We compare our Stage 3 student model with state-of-the-art baselines. Our model surpasses all competitors, including the five teacher models and two recent LMM-based approaches, VQA2 (Jia et al., 2024) and VQAThinker (Cao et al., 2025). Notably, VQA2 is trained on over 157k labeled samples, while VQAThinker leverages reinforcement learning with advanced LMM backbones. In contrast, our weak-to-strong learning strategy achieves state-of-the-art performance without any human-labeled data, underscoring its effectiveness and practical value.
5 Discussion
Developing generalized VQA models remains a fundamental challenge due to the vast diversity of real-world distortions and the strong influence of video content. Supervised learning on human-labeled data cannot feasibly cover this space, highlighting the urgent need for unsupervised and weakly supervised paradigms. In this work, we demonstrate that it is possible to learn from weak VQA models and even surpass their performance. Building on this insight, we propose a framework that integrates diverse homogeneous and heterogeneous VQA teachers through a learning-to-rank formulation, and further enhances generalization via an iterative W2S training strategy, where progressively stronger students are recycled as new teachers. This design enables cumulative transfer of knowledge beyond any single teacher and drives the model’s self-evolution toward increasingly generalized quality assessment.
Looking forward, this paradigm suggests a pathway toward scalable VQA foundation models. The community can leverage a broad spectrum of supervision sources, leveraging expert-domain VQA models (e.g., VMAF for video compression), utilizing powerful LMMs with carefully designed prompt engineering, and employing text-to-video generation algorithms to synthesize videos of varying quality through specified prompts. By unifying these heterogeneous signals, future research may move toward constructing foundation models for VQA that generalize across content domains, distortion types, and application scenarios—ultimately serving as universal quality assessors for both natural and generative videos.
@@ 6 Conclusion
This paper introduces a weak-to-strong (W2S) paradigm for video quality assessment that leverages multiple weak teachers and iterative self-teaching to train stronger students without relying on human annotations. Through the integration of homogeneous and heterogeneous teachers under a ranking-based formulation, and the use of iterative W2S training, our approach consistently surpasses the teacher models across ten benchmarks, with particularly strong gains on challenging out-of-distribution benchmarks. The results highlight the potential of W2S as a scalable and effective alternative to traditional annotation-dependent training pipelines.
LLM Usage Statement. Large language models are used to aid in polishing the writing of this paper, but they are not involved in the research design, experimental process, or analysis.
Ethics Statement. All videos used in this work are obtained through a filtering pipeline that ensures only publicly available content with permissive licenses is included.
Reproducibility Statement. Detailed descriptions of the data processing pipeline, training and inference configurations are provided in the main paper and appendix. Our code link: https://github.com/clh124/W2S-VQA.
Appendix A More Details of Our $D_{\text{w2s}}$ Database
@@ A.1 Analysis of the Collected Videos
As shown in Fig. 8, our dataset is collected from multiple popular social media platforms with relatively uniform sampling, comprising $20\%$ from Bilibili, $20\%$ from Youku, $25\%$ from YouTube, and $35\%$ from TikTok. All videos are obtained through a filtering pipeline that ensures only publicly available content with permissive licenses is included. Notably, our dataset covers a diverse range of content categories, exceeding twenty in total. In addition to common categories such as lifestyle, food, and animals, it also includes specialized categories such as gaming, AI-generated content, and high-resolution content. To illustrate the diversity of our dataset, we present a variety of video samples in Fig. 7, showcasing the broad range of content available in our large-scale video quality assessment (VQA) dataset. Unlike existing datasets, which often focus on specific formats, our dataset encompasses a wider variety of formats, including both landscape and portrait orientations, as well as various resolutions. This diversity enhances the comprehensiveness of our dataset, making it more suitable for evaluating video quality across a wide kinds of scenarios. A detailed breakdown of our database, including pair types and the corresponding number of videos, is provided in Table 2.
@@
Table 2: Statistics of raw videos and video pairs in the $D_{\text{w2s}}$ dataset.
Category Subtype Videos Video Pairs
$D_{\text{w2s}}^{(1)}$ $D_{\text{w2s}}^{(2)}$ $D_{\text{w2s}}^{(3)}$ $D_{\text{w2s}}^{(1)}$ $D_{\text{w2s}}^{(2)}$ $D_{\text{w2s}}^{(3)}$
Ensembling homogeneous teachers - 200k 100k 50k 250k 85k 85k
Integrating heterogeneous teachers Spatial 50k 2k 2k 160k 5k 5k
Temporal 20k 1k 1k 40k 5k 5k
Compression 10k 1k 1k 50k 5k 5k
Total 280k 384k 438k 500k 600k 700k